Slove the quadric equation by completing the square. x^2 -5=8x
Found 2 solutions by solver91311, stanbon:
Answer by solver91311(24713)



Put the first degree term on the left and the constant on the right:



Divide the coefficient on the first degree term by 2, square the result, then add that result to both sides of the equation:



Factor the perfect square on the left and collect terms on the right:



Take the square root of both sides remembering to consider both the positive and negative root.



Add the additive inverse of the constant on the left to both sides:



Done.

John


Answer by stanbon(75887)

Solve the quadratic equation by completing the square.
x^2 -5=8x
----
x^2 - 8x + ? = 5 + ?
----
Complete the square:
x^2 - 8x + 16 = 5 + 16
(x-4)^2 = 21
---
x-4 = +-sqrt(21)
x = [4 +-sqrt(21)]
=======================
Cheers,
Stan H.